rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#1732 — FS#5713 — Robots update

Attached to Project— VPS
Modernization
Infrastructure
CLOSED
100%
We will update the system of upgrade/downgrade of VPS features.
This is going to impact deliveries, changes in hourly and monthly basis.

As a precaution, we will stop robots in connection with these actions tomorrow Wednesday 24th of August 2011 around 8:30am.
The update should not exceed thirty minutes, checking and recovery, including robots.
The service will be on again around 9:00 am.

All actions undertaken in this time period will be queued and played back by the end of the intervention.

Date:  Wednesday, 24 August 2011, 12:48PM
Reason for closing:  Done
Additional comments about closing:  The robots were relaunched. No detected problem. The updating was successful.
Comment by OVH - Wednesday, 24 August 2011, 12:45PM

We will stop the robots and will start again the updating.


Comment by OVH - Wednesday, 24 August 2011, 12:47PM

The updating was finalised, we will relaunch the robots.